Given the quadratic function y = 1 + 4x - x2 . Suppose the domain is restricted to the interval (-½, 4), what is the corresponding range of the function in interval notation?

A. (-5/4, 1)
B. (-5/4, 5)
C. [-5/4, 5)
D. (-5/4, 5]
E. (-5/4, 1]
My solution was to sub both domains in the equation and the answer was -5/4 and 1 but i am confused which one i am going to answer A or E

Answer by Alan3354(69443)

Given the quadratic function y = 1 + 4x - x2 . Suppose the domain is restricted to the interval (-½, 4), what is the corresponding range of the function in interval notation?
----------
f(-1/2) = -5/4
f(4) = 1
----------------
But the function has a maximum of 5 at x=2.
f(2) = 5
--> range of -5/4 to + 5
B